# Unsupervised cell segmentation by fast Gaussian Processes

### v2

@@ 3.2 Generating binary cell masks
The predictive mean at the interior pixel of a cell object typically has a larger value than the one in the background of a microscopy image. Inspired by the IsoData algorithm ridler1978picture in ImageJ abramoff2004image , we developed an automated and robust way to determine the threshold value for separating objects from background noise using the predictive mean from GP models.
@@
As $\Delta c_{k}(\alpha_{m})$ may not be smooth over $\alpha_{m}$, we compute the predictive mean of $\Delta c_{k}(\alpha_{m})$, denoted by $\Delta c^{*}_{k}(\alpha_{m})$, via a GP model with the default Matérn kernel in Equation (3) by the $\sf{RobustGaSP}$ package gu2018robustgasp .
@@
for $k=1,...,\tilde{K}$. We complete panel (C) in Figure 1 by generating binary cell masks for all sub-images. To address a rare scenario with a higher-than-expected object count, which only happened once for all sub-images, we implemented a re-thresholding step discussed in Supplementary Section S3. Finally, each object group is separated and detected via flood fill bhawnesh2020flood , as described in Section S4 in the Supplementary Material.
@@ 3.3 Labeling cell masks by the watershed algorithm
After obtaining the binary image, we apply the watershed algorithm to separate cell objects from $\sf{EBImage}$ package in $\sf{R}$ pau2010ebimage , which is particularly useful for segmenting objects connecting to each other. The watershed algorithm creates a topological representation of an image based on a distance map, which is derived from the binary image representing foreground objects (cells). The distance map assigns a value to each foreground pixel based on its Euclidean distance to the nearest background pixel; higher values represent larger distances from the background. These distance values act as heights in a topological landscape, providing a basis for the watershed function to simulate water through. The negative distance map is then calculated by taking the negative of the distance map. Examples of the binary image and negative distance map are provided in Section S5 in the Supplementary Material.
The local minima of the negative distance map generally correspond to the centers of cell objects and are used as the starting point for the “flooding” process in the watershed algorithm. The water fills the cell object from these local minima as the negative distance map is dunked into water, so the cell object acts as a catch basin as the map floods. As water from different starting points meets, watershed lines are formed, which delineate boundaries for distinct binary cell objects. Additionally, whenever the water from a catch basin meets a background pixel, a watershed line is formed Vincent1991 . These watershed lines separate and create distinct cell objects, and these cell objects can be assigned unique labels.
@@
At the water level 0, the objects are completely segmented, shown in Figure 4 (D). The watershed operation then halts, as this height is used as a stopping point since all cell objects have negative heights in the negative distance map. Additional details of the watershed algorithm, including definitions and edge cases, can be found in Vincent1991 .
@@ 4 Numerical studies of image denoising
We study two examples. In the first example, we studied two scenarios, where the means of the observations follow the Branin function picheny2013benchmark and the linear diffusion equation lu2020prediction , respectively. In the second example, the means of the observations are images of cell nuclei and whole cells.
@@
    where $f(x,t)$ represents the concentration of the diffusing material at location $x$ and time $t$, and $D$ is the diffusion coefficient. We set $D=1$ and discretize the spatial domain $[0,1]$ into $200$ equally spaced grid points. The initial condition is set as $f(x,0)=0$, with a boundary condition at one end, maintaining a constant external concentration of 1. The signal is simulated over the time interval $t\in[0,0.2]$ using $200$ time steps, computed with a numerical solver (soetaert2010solving, ). The observations are generated by $y(x,t)=f(x,t)+\epsilon$, where $\epsilon$ is an independent Gaussian noise. We evaluate the model performance under three configuration with noise variances $\sigma_{0}^{2}\in\{0.05^{2},0.1^{2},0.3^{2}\}$.
@@
We consider two noisy cell microscopy images luo2023molecular : one for cell nuclei and the other for the whole cell. Independent Gaussian noise with $\sigma_{0}^{2}\in\{0.1^{2},0.3^{2},0.5^{2}\}$ is added to each image to generate the observations.
To compare the model performance in noising filtering, we consider the error of the predictive mean of the noisy observations, quantified by the root mean squared error (RMSE):
@@
Figure 5 presents the RMSE in estimating the mean, based on the data-generating processes described in Example 1. Our fast GPs with the Matérn kernel in Equation (3) consistently achieve the highest accuracy across varying noise levels in both scenarios. This is because GPs with separable kernels capture spatial dependencies in both input directions, which are crucial for modeling images. In contrast, PCA implicitly assumes that the prior in one direction is independent, as the corresponding probabilistic model assumes that the latent factors are distributed as a standard Gaussian distribution tipping1999probabilistic . Though DMD and FMOU assume the latent factors are from Gaussian processes, which capture output correlation over one input dimension, the latent factor loading matrix is estimated without modeling the correlation over the other input by a kernel function. In comparison, Fast GPs with Matérn kernel directly model the correlation over two inputs through a product kernel, which induces better predictions of the signals. Additionally, since DMD implicitly assumes a noise-free model gu2024probabilistic , leading to degraded performance as noise variance increases. Figure S6 in the Supplementary Material shows the predictive signal by Fast GPs with Matérn kernel from noisy observations generated by the Branin function and the linear diffusion equation, which demonstrates close alignment with the ground truth.
@@
Figure 6 compares the accuracy of five approaches in denoising cell nuclei and whole-cell images in Example 2. The fast GPs with Matérn and exponential kernels significantly outperform PCA, FMOU and DMD across all settings, owing to their ability to estimate correlation parameters in both directions. Figure 7 plots the signal, noisy observations, and predictive mean from the fast GPs with a Matérn kernel in Equation (3) for denosing the noisy cell images. The predictive mean is close to the latent mean of the observations.
@@ 5.1 Criteria and methods
We use microscopy images of human dermal fibroblasts (hDFs) in luo2023molecular for testing different approaches. More details on the experimental conditions and techniques can be found in Section S7 in the Supplementary Material. These images contain both nuclei and cytoplasm channels, and evaluation of the proposed GP-based segmentation method was performed on both channels.
To evaluate the accuracy of the truth and segmentation results, we compare the area of overlap between the detected object and the true object relative to the overall union area of the two objects Rezatofighi_2019_CVPR . Let $g$ represent a ground truth mask and $p$ a predicted mask. The Intersection over Union (IoU) metric between $g$ and $p$ is defined as:
@@
The average precision metric ($\text{AP}(\alpha)$) at each IoU threshold $\alpha\in[0,1]$ is often used for evaluating the correction of image segmentation (stringer2021cellpose, ). For $\alpha\in[0,1]$, a True Positive (TP) means that the ground truth mask $g$ is matched to a predicted mask $p$ with $\text{IoU}(g,p)\geq\alpha$. Otherwise, unmatched predicted masks are counted as False Positives (FP), and unmatched ground truth masks are counted as False Negatives (FN). The $\mbox{AP}(\alpha)$ is calculated as follows:
@@
We compare ImageJ segmentation and the GP-based unsupervised detection algorithm developed in this work. We followed the conventional processing steps in ImageJ to generate the cell masks schneider2012nih . All input images are converted to grayscale. To detect the foreground and background pixels, a threshold was set for each image using the default methods. To reduce noises in ImageJ, a despeckle operation, which applies a $3\times 3$ median filter, was then applied to each image. A watershed algorithm was used to ensure better separation of connecting cells. Finally, the cell masks were generated. The cell size is selected to be larger than $30$ pixels. Then, the binary image and the labeled mask can be generated in ImageJ. A detailed description of each step in this process for ImageJ can be found in Section S8 in the Supplementary Material.
@@ 5.2 Image segmentation results for cell nuclei
In panel (A) of Figure 9, we plot the $\mbox{AP}(\alpha)$ between the GP-based segmentation and ImageJ segmentation methods by solid lines and dashed lines, respectively, for 5 test images of nuclei at distinct threshold $\alpha$. For any image and threshold level, the solid line is almost always above the dashed line with the same color, as the GP-based method consistently achieves higher AP than ImageJ at any threshold level $\alpha$. The performance gap becomes more obvious at intermediate IoU thresholds between $0.6$ to $0.7$, which means that the GP-based method is better at recovering the overall structure of the object than ImageJ. The AP score of both methods drops at a high IoU threshold, as the annotated object may not be fully accurate, which can affect defining the truth. Panel (B) in Figure 9 shows the distribution of these AP scores from GPs and ImageJ segmentation methods averaged at each threshold. Similar to the results in panel (A), the GP-based method achieves higher median AP scores across all thresholds. The difference between the GP-based method and ImageJ is more pronounced at a high threshold level, such as $\alpha=0.75$ or $\alpha=0.8$, where the boxes of two colors nearly have no overlap, indicating that GP-based method better capture the fine-scale boundary details than ImageJ.
@@ 6 Concluding remarks
In this study, we have presented a scalable cell image segmentation method for both the cell nucleus and cytoplasm. By leveraging fast Gaussian processes, automated thresholding and watershed operations, we shown the process is not only more automated, but the results are more precise than other unsupervised segmentation methods, such as ImageJ. A key strength of the proposed segmentation algorithm is that it does not rely on parameter tuning or labeled training datasets. The results of our unsupervised method demonstrate versatility across different cell channels, making it an appealing option for more general segmentation tasks beyond cell segmentation.
Several future directions are worth exploring. First, modern microscopes provide time-lapse videos for tracking the changes in cellular behavior over time. A future goal is to integrate our segmentation methods with particle tracking and linking algorithms that solve a linear assignment problem jaqaman2008robust . As our approach does not restrict the shape of the object, it can be extended for image segmentation with objects undergoing dynamical changes, including microscopy images for the nucleation and growth process in crystallization wu2022nucleation and protein dynamics lee2021myosin . Second, when labeled image data are available, it is worthwhile to explore how these data can be used for constraining the shapes of the objects to improve estimation and simultaneously enable segmenting new objects not in the database. Third, the majority of cell segmentation methods are developed for 2D microscopy images. It is of interest to develop cell segmentation methods for microscopy that capture cell behaviors in a 3D environment, such as the orientational order and alignment huang2025cell . Similar to some popular cell segmentation tools, such as Cellpose stringer2021cellpose , our approach aims to segment images that contain the same type of cell. A future direction is to leverage techniques in tools such as CellProfiler carpenter2006cellprofiler to segment images that contain multiple types of cells. Furthermore, the uncertainty is often not quantified by the image segmentation methods, as the segmentation process often contains multiple steps, and the uncertainty from the convolutional neural network is an open question. As we model the image data with a probabilistic model, a future goal is to propagate the uncertainty throughout the analysis for uncertainty quantification.
@@ Acknowledgement
The $t=i+(j-1)n_{1}$ diagonal term of $\hat{\sigma}^{2}(\mathbf{R}-\mathbf{R}_{2}\otimes\mathbf{R}_{1})$ is the predictive variance at pixel location $(i,j)$, which follows
@@
$=$ $\hat{\sigma}^{2}\left(1-\sum^{n_{2}}_{j^{\prime}=1}\sum^{n_{1}}_{i^{\prime}=1}\frac{\tilde{r}^{2}_{i,i^{\prime},1}\tilde{r}^{2}_{j,j^{\prime},2}}{\lambda_{i,1}\lambda_{j,2}+\eta}\right),$
where $\tilde{r}_{i,i^{\prime},1}$ is the $i^{\prime}$th term of the vector $\mathbf{r}^{T}_{i,1}\mathbf{U}_{1}$ and $\tilde{r}_{j,j^{\prime},2}$ is the $j^{\prime}$th term of the vector $\mathbf{r}^{T}_{j,2}\mathbf{U}_{2}$, for $i^{\prime}=1,...,n_{1}$ and $j^{\prime}=1,...,n_{2}$, with $\mathbf{r}^{T}_{i,1}=(K_{1}(x_{i,1},x_{1,1}),...,K_{1}(x_{i,1},x_{n_{1},1}))^{T}$ and $\mathbf{r}^{T}_{j,2}=(K_{2}(x_{j,2},x_{1,2}),...,K_{1}(x_{j,2},x_{n_{2},2}))^{T}$.

### v3

@@ 3.2 Generating binary cell masks
The predictive mean at the interior pixel of a cell object typically has a larger value than the one in the background of a microscopy image. Inspired by the IsoData algorithm Ridler et al. (1978) in ImageJ Abràmoff et al. (2004), we developed an automated and robust way to determine the threshold value for separating objects from background noise using the predictive mean from GP models.
@@
As $\Delta c_{k}(\alpha_{m})$ may not be smooth over $\alpha_{m}$, we compute the predictive mean of $\Delta c_{k}(\alpha_{m})$, denoted by $\Delta c^{*}_{k}(\alpha_{m})$, via a GP model with the default Matérn kernel in Equation (3) by the $\sf{RobustGaSP}$ package Gu et al. (2019).
@@
for $k=1,...,\tilde{K}$. We complete panel (C) in Figure 1 by generating binary cell masks for all sub-images. To address a rare scenario with a higher-than-expected object count, which only happened once for all sub-images, we implemented a re-thresholding step discussed in Section S3 in the Supplementary Material. Finally, each object group is separated and detected via flood fill Bhawnesh et al. (2020), as described in Section S4 in the Supplementary Material.
@@ 3.3 Labeling cell masks by the watershed algorithm
After obtaining the binary image, we apply the watershed algorithm to separate cell objects from $\sf{EBImage}$ package in $\sf{R}$ Pau et al. (2010), which is particularly useful for segmenting objects connecting to each other. The watershed algorithm creates a topological representation of an image based on a distance map, which is derived from the binary image representing foreground objects (cells). The distance map assigns a value to each foreground pixel based on its Euclidean distance to the nearest background pixel; higher values represent larger distances from the background. These distance values act as heights in a topological landscape, providing a basis for the watershed function to simulate water through. The negative distance map is then calculated by taking the negative of the distance map. Examples of the binary image and negative distance map are provided in Section S5 in the Supplementary Material.
The local minima of the negative distance map generally correspond to the centers of cell objects and are used as the starting point for the “flooding” process in the watershed algorithm. The water fills the cell object from these local minima as the negative distance map is dunked into water, so the cell object acts as a catch basin as the map floods. As water from different starting points meets, watershed lines are formed, which delineate boundaries for distinct binary cell objects. Additionally, whenever the water from a catch basin meets a background pixel, a watershed line is formed Vincent and Soille (1991). These watershed lines separate and create distinct cell objects, and these cell objects can be assigned unique labels.
@@
At the water level 0, the objects are completely segmented, shown in Figure 4 (D). The watershed operation then halts, as this height is used as a stopping point since all cell objects have negative heights in the negative distance map. Additional details of the watershed algorithm, including definitions and edge cases, can be found in Vincent and Soille (1991).
@@ 4 Numerical studies of image denoising
We study two examples. In the first example, we studied two scenarios, where the means of the observations follow the Branin function Picheny et al. (2013) and the linear diffusion equation Lu and Tartakovsky (2020), respectively. In the second example, the means of the observations are images of cell nuclei and whole cells.
@@
    where $f(x,t)$ represents the concentration of the diffusing material at location $x$ and time $t$, and $D$ is the diffusion coefficient. We set $D=1$ and discretize the spatial domain $[0,1]$ into $200$ equally spaced grid points. The initial condition is set as $f(x,0)=0$, with a boundary condition at one end, maintaining a constant external concentration of 1. The signal is simulated over the time interval $t\in[0,0.2]$ using $200$ time steps, computed with a numerical solver (Soetaert et al., 2010). The observations are generated by $y(x,t)=f(x,t)+\epsilon$, where $\epsilon$ is an independent Gaussian noise. We evaluate the model performance under three configurations with noise variances $\sigma_{0}^{2}\in\{0.05^{2},0.1^{2},0.3^{2}\}$.
@@
We consider two noisy cell microscopy images Luo et al. (2023): one for cell nuclei and the other for the whole cell. Independent Gaussian noise with $\sigma_{0}^{2}\in\{0.1^{2},0.3^{2},0.5^{2}\}$ is added to each image to generate the observations.
To compare the model performance in noise filtering, we consider the error of the predictive mean of the noisy observations, quantified by the root mean squared error (RMSE):
@@
Figure 5 presents the RMSE in estimating the mean, based on the data-generating processes described in Example 1. Our fast GPs with the Matérn kernel in Equation (3) consistently achieve the highest accuracy across varying noise levels in both scenarios. This is because GPs with separable kernels capture spatial dependencies in both input directions, which are crucial for modeling images. In contrast, PCA implicitly assumes that the prior in one direction is independent, as the corresponding probabilistic model assumes that the latent factors are distributed as a standard Gaussian distribution Tipping and Bishop (1999). Though DMD and FMOU assume the latent factors are from Gaussian processes, which capture output correlation over one input dimension, the latent factor loading matrix is estimated without modeling the correlation over the other input by a kernel function. In comparison, Fast GPs with Matérn kernel directly model the correlation over two inputs through a product kernel, which induces better predictions of the signals. Additionally, since DMD implicitly assumes a noise-free model Gu et al. (2024), leading to degraded performance as noise variance increases. Figure S6 in the Supplementary Material shows the predictive signal by Fast GPs with Matérn kernel from noisy observations generated by the Branin function and the linear diffusion equation, which demonstrates close alignment with the ground truth.
@@
Figure 6 compares the accuracy of five approaches in denoising cell nuclei and whole-cell images in Example 2. The fast GPs with Matérn and exponential kernels significantly outperform PCA, FMOU and DMD across all settings, owing to their ability to estimate correlation parameters in both directions. Figure 7 plots the signal, noisy observations, and predictive mean from the fast GPs with a Matérn kernel in Equation (3) for denoising the noisy cell images. The predictive mean is close to the latent mean of the observations.
@@ 5.1 Criteria and methods
We use microscopy images of human dermal fibroblasts (hDFs) in Luo et al. (2023) for testing different approaches. More details on the experimental conditions and techniques can be found in Section S7 in the Supplementary Material. These images contain both nuclei and cytoplasm channels, and evaluation of the proposed GP-based segmentation method was performed on both channels.
To evaluate the accuracy of the truth and segmentation results, we compare the area of overlap between the detected object and the true object relative to the overall union area of the two objects Rezatofighi et al. (2019). Let $g$ represent a ground truth mask and $p$ a predicted mask. The Intersection over Union (IoU) metric between $g$ and $p$ is defined as:
@@
The average precision metric ($\text{AP}(\alpha)$) at each IoU threshold $\alpha\in[0,1]$ is often used for evaluating the correction of image segmentation (Stringer et al., 2021). For $\alpha\in[0,1]$, a True Positive (TP) means that the ground truth mask $g$ is matched to a predicted mask $p$ with $\text{IoU}(g,p)\geq\alpha$. Otherwise, unmatched predicted masks are counted as False Positives (FP), and unmatched ground truth masks are counted as False Negatives (FN). The $\mbox{AP}(\alpha)$ is calculated as follows:
@@
We compare ImageJ segmentation and the GP-based unsupervised detection algorithm developed in this work. We followed the conventional processing steps in ImageJ to generate the cell masks Schneider et al. (2012). All input images are converted to grayscale. To detect the foreground and background pixels, a threshold was set for each image using the default methods. To reduce noises in ImageJ, a despeckle operation, which applies a $3\times 3$ median filter, was then applied to each image. A watershed algorithm was used to ensure better separation of connecting cells. Finally, the cell masks were generated. The cell size is selected to be larger than $30$ pixels. Then, the binary image and the labeled mask can be generated in ImageJ. A detailed description of each step in this process for ImageJ can be found in Section S8 in the Supplementary Material.
@@ 5.2 Image segmentation results for cell nuclei
In panel (A) of Figure 9, we plot the $\mbox{AP}(\alpha)$ between the GP-based segmentation and ImageJ segmentation methods by solid lines and dashed lines, respectively, for 5 test images of nuclei at distinct threshold $\alpha$. For any image and threshold level, the solid line is almost always above the dashed line with the same color, as the GP-based method consistently achieves higher AP than ImageJ at any threshold level $\alpha$. The performance gap becomes more obvious at intermediate IoU thresholds between $0.6$ to $0.7$, which means that the GP-based method performs better at recovering the overall structure of the object than ImageJ. The AP score of both methods drops at a high IoU threshold, as the annotated object may not be fully accurate, which can affect defining the truth. Panel (B) in Figure 9 shows the distribution of these AP scores from GPs and ImageJ segmentation methods averaged at each threshold. Similar to the results in panel (A), the GP-based method achieves higher median AP scores across all thresholds. The difference between the GP-based method and ImageJ is more pronounced at a high threshold level, such as $\alpha=0.75$ or $\alpha=0.8$, where the boxes of two colors nearly have no overlap, indicating that GP-based method better capture the fine-scale boundary details than ImageJ.
@@ 6 Concluding remarks
In this study, we have presented a scalable cell image segmentation method for both the cell nucleus and cytoplasm. By leveraging fast Gaussian processes, automated thresholding and watershed operations, we have shown that the process is not only more automated, but the results are more precise than other unsupervised segmentation methods, such as ImageJ. A key strength of the proposed segmentation algorithm is that it does not rely on parameter tuning or labeled training datasets. The results of our unsupervised method demonstrate versatility across different cell channels, making it an appealing option for more general segmentation tasks beyond cell segmentation.
Several future directions are worth exploring. First, modern microscopes provide time-lapse videos for tracking the changes in cellular behavior over time. A future goal is to integrate our segmentation methods with particle tracking and linking algorithms that solve a linear assignment problem Jaqaman et al. (2008). As our approach does not restrict the shape of the object, it can be extended for image segmentation with objects undergoing dynamical changes, including microscopy images for the nucleation and growth process in crystallization Wu et al. (2022) and protein dynamics Lee et al. (2021). Second, when labeled image data are available, it is worthwhile to explore how these data can be used for constraining the shapes of the objects to improve estimation and simultaneously enable segmenting new objects not in the database. Third, the majority of cell segmentation methods are developed for 2D microscopy images. It is of interest to develop cell segmentation methods for microscopy that capture cell behaviors in a 3D environment, such as the orientational order and alignment Huang et al. (2025). Similar to some popular cell segmentation tools, such as Cellpose Stringer et al. (2021), our approach aims to segment images that contain the same type of cell. A future direction is to leverage techniques in tools such as CellProfiler Carpenter et al. (2006) to segment images that contain multiple types of cells. Furthermore, the uncertainty is often not quantified by the image segmentation methods, as the segmentation process often contains multiple steps, and the uncertainty from the convolutional neural network is an open question. As we model the image data with a probabilistic model, a future goal is to propagate the uncertainty throughout the analysis for uncertainty quantification.
@@ Acknowledgement
The $t=i+(j-1)n_{1}$ diagonal term of $\hat{\sigma}^{2}\mathbf{R}^{*}$ is the predictive variance at pixel location $(i,j)$, which follows
@@
$=$ $\hat{\sigma}^{2}\left(1-\sum^{n_{2}}_{j^{\prime}=1}\sum^{n_{1}}_{i^{\prime}=1}\frac{\tilde{r}^{2}_{i,i^{\prime},1}\tilde{r}^{2}_{j,j^{\prime},2}}{\lambda_{i^{\prime},1}\lambda_{j^{\prime},2}+\eta}\right),$